OpenStudy (anonymous):

Prove that if M is nilpotent, then det(M)=0

OpenStudy (experimentx):

can it be equivalently said that if inverse of matrix exist then it's determinant exist??

OpenStudy (anonymous):

if the inverse of a matrix exists, then det(A) is not zero

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

maybe it has to do with that if at some point det(A^m)=0, then det(A)=0 because you can't multiply by anything else to get 0

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Nilpotent_matrix Let M be non zero matrix. M^k = M x M^(k-1) = 0 let M' (if it exist) be inverse of M, M'M x M^(k-1) = 0 Inductively you can prove that M = 0, which means M' is not inverse of M, so There is no inverse of M.

OpenStudy (anonymous):

thanks

OpenStudy (experimentx):

seriously ,,, i don't know anything about this subject.

OpenStudy (anonymous):

no? it's advanced linear algebra....I used the fact that det(AB)=det(A)det(B)

OpenStudy (experimentx):

oh ... thats great.